Fifteen people are randomly assigned, so that 5 people exercise in the morning, afternoon, or night. After 8 weeks, their weight loss is measured in pounds. Compute SSwithin.SS morning (EXmorning) morning n (EX afternoon) SS afternoon afternoon n (EXnight) SS night night n SS within(error) = SS mor

A. 3,719.2

B. 22.8

C. 3,615.6

D. 525.6

Can you please show me how you get your answer? I'm writing it out and I keep getting 525.6 but I keep being told that it's wrong. Thank you!
Answer #1

Let the \sum = SUM

SS morning = SUM(X2 morning) - SUM(X morning)2/n = 319 - (39)2/5 = 319 - (1521 / 5) = 319 - 304.2 = 14.8

SS afternoon = SUM(X2 afternoon) - SUM(X afternoon)2/n = 307 - (39)2/5 = 307 - (1521 / 5) = 307 - 304.2 = 2.8

SS night = SUM(X2 night) - SUM(X night)2/n = 151 - (27)2/5 = 151 - (729 / 5) = 151 - 145.8 = 5.2

Therefore SS within = SS morning + S afternoon + SS night = 14.8 + 2.8 + 5.2 = 22.8

Option B